OpenStudy (anonymous):

What are the solutions of 3x^2 + 6x + 6 = 0? a. -2 ± i b. -2 ± 2i c. -1 ± i d. 1 ± i

OpenStudy (agent0smith):

Do you know how to use the quadratic formula? for \[\large ax^2 + bx + c=0\] \[\Large x= \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (agent0smith):

@lowcard2 you'll learn a lot more if you get involved as opposed to just reading answers.

OpenStudy (anonymous):